# Few-shot Object Localization

### v1

@@ Abstract
Existing few-shot object counting tasks primarily focus on quantifying the number of objects in an image, neglecting precise positional information. To bridge this research gap, this paper introduces the novel task of Few-Shot Object Localization (FSOL), which aims to provide accurate object positional information. This task achieves generalized object localization by leveraging a small number of labeled support samples to query the positional information of objects within corresponding images. To advance this research field, we propose an innovative high-performance baseline model. Our model integrates a dual-path feature augmentation module to enhance shape association and gradient differences between supports and query images, alongside a self-query module designed to explore the association between feature maps and query images. Experimental results demonstrate a significant performance improvement of our approach in the FSOL task, establishing an efficient benchmark for further research.
@@
Object localization; Few-shot object localization; Few-shot object counting; Few-shot learning.
@@ I Introduction
Object localization [1, 2, 3], a fundamental task in computer vision, has witnessed remarkable progress, propelled by the advent of deep learning techniques. The precise object localization within images holds critical significance across various applications, including autonomous vehicles [4], surveillance systems [5], medical image analysis [6, 7], and crowd management [8]. Despite the strides made in this field, existing methods predominantly rely on the availability of abundant labeled data to train models with high accuracy. However, the acquisition of such labeled datasets in real-world scenarios often poses formidable challenges, owing to the associated expenses and time constraints. In response to this challenges, few-shot learning emerges as a promising paradigm to alleviate the dependence on extensive labeled datasets [9, 10, 11]. By empowering models to learn knowledge from a little number of labeled examples, few-shot learning presents a compelling way to enhance generalization capabilities. This method proves particularly advantageous in scenarios where the availability of substantial labeled data becomes impractical or unfeasible.
@@
Nevertheless, the integration of few-shot learning and object localization remains unexplored. To bridge this research gap, this paper first introduces the Few-Shot Object Localization (FSOL) task. In contrast to few-shot object counting tasks [12] that primarily focus on quantity analysis, FSOL emphasizes not only recognizing objects but also providing accurate positional information within the image. By focusing on a limited number of labeled support samples, FSOL aims to achieve a generalized object localization capability across diverse scenarios. As shown in Fig. 1, during the training phase, the model learns from labeled support samples of known categories. Subsequently, during the testing phase, the model demonstrates a remarkable capacity to extend its knowledge to entirely new categories, significantly enhancing its overall generalization and adaptability.
@@
Figure 1: Demonstration of the Few-Shot Object Localization (FSOL) task. During the training phase, the model predicts the location map using one given support sample and corresponding query image. It then optimizes itself by calculating the loss between the ground truth and the predicted location map. In the testing phase, the trained model predicts the location map of novel class samples on corresponding query images, which did not appear in the training phase.
@@
This study aims to advance the research field by introducing a high-performance benchmark model for the FSOL task. Within this task, we have identified two primary challenges: 1) Appearance gap between intra-class objects: The significant intra-class variations among objects within the query image create an appearance gap compared to samples in the support image, thereby impacting the query’s accuracy, as shown in Fig. 2(a). 2) Object omission due to inter-object occlusion: the model struggles to accurately distinguish the dense, overlapping objects in the query image, leading to a decrease in localization recall rate, as shown in Fig. 2(b).
@@
Figure 2: Difficulties in few-shot object localization which causing negative influence: a) Appearance gap between intra-class objects; b) Object omission due to inter-object occlusion.
To address the two key challenges in FSOL task, we propose an innovative dual-path feature augmentation module. Firstly, to mitigate the significant intra-class variations of objects in the query image, we introduce a deformable convolution branch [13, 14]. This branch enables the model to effectively learn the variations of intra-class features, thereby enhancing localization accuracy by capturing deformable features present in both support samples and query objects. Secondly, to tackle the frequent occurrence of object omission, we introduce a cross-center difference convolution branch [15]. This branch focuses on capturing gradient similarity in support samples and query objects, significantly improving the model’s discriminability of object features and reducing the occurrence of object omission. To optimize the model’s efficiency and generalization, we implement weight-sharing between support and query images, reducing the number of parameters and bridging the gap between support samples and query objects. Finally, to comprehensively capture the intrinsic structure, texture, and patterns of images, we introduce 3D convolution, considering the association of channel information, thereby enhancing the representational capability of features and the overall performance of the model. Finally, the query image performs convolution with the support image in three dimensions to obtain a similarity map reflecting the object location information.
Further, leveraging the original query image to enhance the obtained similarity map has emerged as a promising strategy. Current methods commonly involve directly adding the query image to the similarity feature map, employing a residual connection technique to preserve object information from the original image and thereby optimizing the similarity feature map [16]. However, this direct addition strategy introduces considerable noise from the query image, making it unsuitable for localization tasks that demand high accuracy. Therefore, inspired by the self-support matching [17], we utilize the similarity matrix calculated between the query image and similarity map for weighting, aiming to introduce the query image’s information more accurately while mitigating excessive noise. This innovative design is expected to further improve the accuracy and robustness of the model, opening up new directions for the research of FSOL task.
In this paper, we introduces the pioneering task of FSOL and presents an innovative high-performance baseline model. To tackle the challenges posed by significant intra-class variations and occlusions in localization tasks, we propose a dual-path feature augmentation (DFA) module aimed at enhancing shape association and gradient differences between support and query instances. Moreover, to effectively leverage query images for enhancing similarity maps, we introduce a self-query (SQ) module to explore intricate associations between feature maps and query images. Experimental results demonstrate the substantial performance improvement of our approach in the FSOL task, establishing an efficient benchmark for further research in object localization under limited data scenarios. In summary, the contributions of this paper can be outlined as follows:
  • We first propose the task of few-shot object localization, providing a new research direction for object localization in scenarios with limited labeled data, and establish a high-performance benchmark.
  • Innovatively, we adopt the dual-path feature augmentation module, paying attention to the relationship between channels and image features simultaneously, significantly improving the performance of the task.
  • We introduce a self-query module, cleverly uses the similarity matrix for weighting, solving the problem of object omission, and effectively improving the accuracy of localization.
@@ II Related Work
This paper is the first work to propose the task of few-shot object localization. Therefore, in this section, we briefly review the research in other fields related to this paper. First, we outline some classic works for the object localization task, then briefly introduce the development of few-shot learning, and finally introduce the currently prevalent works for few-shot object counting.
@@ II-A Object Localization
Object localization aims to precisely determine the position of objects in images or videos. Unlike object detection, which involves bounding box information, object localization focuses solely on object position. This localization process typically involves two stages: generating a location map through model prediction, followed by post-processing to extract object position details [18, 19, 20]. The evolution from density maps [21, 22, 23] to location maps [24, 18, 19] emphasizes accurate object localization and high-quality gradient data for model training. Traditional density maps struggle with precise object localization, prompting innovations like the IDT map [24], designed to provide accurate object positions through distance flipping. Despite improvements, IDT maps often contain significant background noise. Addressing this, Liang et al. [18] introduced the Focal Inverse Distance Transform (FIDT) map, offering enhanced gradient information. However, issues persist, such as rapid foreground decay and slow background decline. To mitigate these, Li et al. [19] proposed using exponential functions, moderating foreground decay and expediting background zeroing, resulting in more optimal location maps.
@@ II-B Few-shot Learning
Few-shot Learning methods offer competitive performance even with limited annotated data, a critical attribute in today’s era of data scarcity [32]. This paper proposes integrating few-shot Learning techniques with object localization tasks to explore innovative approaches for enhancing object localization performance under few-shot scenarios.
@@ II-C Few-shot Object Counting
The Few-shot Object Counting task aims to describe object categories using support images and then count the instances of these categories in the query image. Previous research [33, 34] primarily focused on enumerating specific objects within single categories, constraining the task’s generalization capability. Lu et al. [35] introduced class-agnostic counting, developing a GMN model capable of counting objects across any category by reframing the problem as an image self-similarity matching task. In contrast, Yang et al. [36] employed multi-scale computation to assess similarity between support and query images, transforming object counting into a similarity matching challenge. Due to the absence of dedicated few-shot object counting datasets, Ranjan et al. [37] introduced the FSC-147 dataset, which includes 147 object categories and 6000 images, and proposed the FamNet model based on similarity comparison techniques. More recently, You et al. [16] integrated feature enhancement and similarity matching concepts, achieving state-of-the-art performance in few-shot object counting. These studies contribute both theoretically and empirically to advancing the few-shot object counting task.
@@ III Method
This study proposes a novel Few-Shot Object Localization (FSOL) task to enhance object localization accuracy and generalization in few-shot scenarios. To this end, we employ three key strategies, as shown in Fig. 3. Firstly, our method emphasizes capturing deformation and gradient information within the image to better understand the association between query and support features. Secondly, we integrate 3D convolution to comprehensively consider various dimensions’ impact on feature similarity comparison. Lastly, we introduce a module utilizing original query features to guide similarity distribution, allowing the model to utilize optimized similarity information for identifying and locating challenging samples. This reduces object omissions and enhances the model’s generalization.
Figure 3: Demonstration of our proposed FSOL Pipeline. Given the query and support images, the query feature $F_{Q}$ is extracted from the query image and the support feature $F_{S}$ is cropped from $F_{Q}$. The Dual-path Feature Augmentation (DFA) module first enhances deformation and gradient information in $F_{Q}$ and $F_{S}$ through deformation and gradient branches, outputting the deformation-enhanced $F_{Q}^{D}$, $F_{S}^{D}$ as well as gradient-enhanced $F_{Q}^{C}$ and $F_{S}^{C}$. Then, DFA uses 3D convolution to fuse the stacked $F_{Q}^{D}$ and $F_{S}^{D}$ using stacked $F_{Q}^{C}$ and $F_{S}^{C}$ as convolution kernel weights, obtaining the similarity map $S$ between query image and support image. Then, the Self-query (SQ) module accepts $S$ as input and uses the original $F_{Q}$ to guide the object’s distribution information in $S$, subsequently outputting the optimized similarity map $S_{SQ}$. The $S_{SQ}$ then be sent to regression head to get the final location map.
@@ III-A Dual-path Feature Augmentation
In this study, we introduce a Dual-path Feature Augmentation (DFA) module aimed at mitigating significant intra-class appearance gaps and addressing object omission challenges in FSOL task. The DFA module comprises two key components: a deformation branch and a gradient branch. The deformation branch facilitates the model in learning flexible matching relationships within intra-class features, thereby improving localization accuracy by capturing deformable features in both support and query images. On the other hand, the gradient branch emphasizes comparing gradient similarities between support and query images during similarity computation. This enhances the model’s ability to discern object features, thereby effectively reducing omissions. The subsequent sections will elaborate on the vanilla 2D convolution, followed by a detailed explanation of the deformation branch and gradient branch.
@@ III-A1 Vanilla 2D Convolution
In convolutional neural networks, the vanilla 2D convolution serves as a fundamental operator, comprising two primary steps: 1) Sampling local neighboring regions on the input feature map $x$ using $N$ sampling points; 2) Aggregating the sampled values through learnable weights $w$. Therefore, for the position $p$ on the input feature map $x$, the corresponding value at the same position $p$ on the output feature map $y$ can be mathematically expressed as:
@@
where $p_{k}$ represents the preset offset of the enumerated position p in the local neighbor area. For example, when $N=9$, there are nine sampling points, $p_{k}=\{(-1,-1),(-1,0),\ldots,(0,1),(1,1)\}$. In this paper, we initially apply vanilla 2D convolution to process the image. For an input image $P\in\mathbb{R}^{1\times 3\times 512\times 512}$, we use the first three stage outputs of ResNet50, resulting in outputs of 256, 512, and 1024 channels, respectively. These outputs are concatenated into a feature tensor $F\in\mathbb{R}^{1\times 1792\times 128\times 128}$. Subsequently, query features $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ are obtained through the aforementioned vanilla 2D convolution. For our primary experiment, the support feature $F_{S}\in\mathbb{R}^{1\times 256\times x\times y}$ is cropped from the query feature $F_{Q}$ based on the dimensions of the original input sample. Following an adaptive pooling layer, the support feature is resized to $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$. Subsequently, these features are forwarded to both the deformation branch and the gradient branch for further processing.
@@ III-A2 Deformation Branch
In FSOL task, the object within the query image often exhibits substantial intra-class variation, distinct from the object depicted in the support image, thereby resulting in inaccurate queries. To address this issue, we aim to enhance the model’s capability to capture deformable information present in both query and support features. This enhancement seeks to improve query accuracy by enabling the model to better model object deformations in both images.
To this end, we incorporate a Deformable Convolution (DC [13, 14]) branch. Unlike traditional convolution, which relies on fixed sampling points within the receptive field, DC introduces an offset to each sampling point, thereby making their positions variable. This offset is generated by another convolution layer. In Fig. 4(a), we depict the disparity between sampling points in vanilla and deformable convolution. While vanilla convolution tends to integrate redundant information into features, deformable convolution diminishes redundancy by iteratively optimizing sampling points. To assess the relevance of the introduced area, DC incorporates an additional weight coefficient. When the area is deemed irrelevant, this coefficient is set to 0. Mathematically, for the same position $p$ in the input feature map $x$ and the output feature map $y$, the relationship between the values $x(p)$ and $y(p)$ after deformable convolution can be expressed as follows:
@@
where $\Delta p_{k}$ represents the offset learned by each sampling point, and $\Delta m_{k}$ represents the weight coefficient in the range [0, 1].
both the original query and support features fed into the deformation branch undergo processing by a shared Deformable Convolution (DC) layer to emphasize their deformation information.
In our framework, both the original query and support features fed into deformation branch undergo processing by a shared DC layer to highlight their deformation information. Specifically, the query feature $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$ each go through a deformable convolution layer, resulting in the enhanced query feature $F_{Q}^{D}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}^{D}\in\mathbb{R}^{1\times 256\times 3\times 3}$. This design not only reduces the amount of computation but also unifies the deformation distribution of both query and support features, alleviating the appearance gap between intra-class samples.
Figure 4: Demonstration of two convolutional strategies for enhancing deformation and gradient information in support and query images: (a) Vanilla convolution vs. Deformable convolution: Vanilla convolution utilizes fixed sampling points, potentially introducing noise, while deformable convolution adjusts sampling points, reducing background noise and improving adaptability. (b) Cross-center difference convolution (CCD-C) process: CCD-C computes differences between neighboring pixels around the central pixel and uses these differences as weights to generate the final output. This approach captures subtle image changes like texture, edges, and fine details.

### v3

@@ Abstract
Existing object localization methods are tailored to locate specific classes of objects, relying heavily on abundant labeled data for model optimization. However, acquiring large amounts of labeled data is challenging in many real-world scenarios, significantly limiting the broader application of localization models. To bridge this research gap, this paper defines a novel task named Few-Shot Object Localization (FSOL), which aims to achieve precise localization with limited samples. This task achieves generalized object localization by leveraging a small number of labeled support samples to query the positional information of objects within corresponding images. To advance this field, we design an innovative high-performance baseline model. This model integrates a dual-path feature augmentation module to enhance shape association and gradient differences between supports and query images, alongside a self query module to explore the association between feature maps and query images. Experimental results demonstrate a significant performance improvement of our approach in the FSOL task, establishing an efficient benchmark for further research. All codes and data are available at https://github.com/Ryh1218/FSOL.
@@
Object localization; Few-shot learning; Few-shot object localization; Object counting.
@@ I Introduction
Object localization [1, 2, 3], a fundamental task in computer vision, has witnessed remarkable progress propelled by deep learning techniques. Precise object localization within images is critically important across various applications, including autonomous vehicles [4], surveillance systems [5], medical image analysis [6, 7], and crowd management [8]. Despite significant advancements, existing methods predominantly rely on abundant labeled data to train high-accuracy models. However, acquiring such labeled datasets in real-world scenarios often poses formidable challenges due to associated expenses and time constraints. In response to these challenges, few-shot learning has emerged as a promising paradigm to reduce dependence on extensive labeled datasets [9, 10, 11]. By enabling models to learn from a limited number of labeled examples, few-shot learning enhances generalization capabilities, proving particularly advantageous in scenarios where substantial labeled data is impractical or unfeasible.
@@
With the advances in few-shot learning, accurate object localization without a large amount of labeled data is achievable yet unexplored. Consequently, we further explore object localization under few-shot settings, namely, when given an image with only a few samples labeled, locating other same-category samples solely with the assistance of the labeled ones. We term this problem the Few-Shot Object Localization (FSOL) task. Unlike few-shot object counting tasks [12] that primarily focus on quantity analysis, FSOL emphasizes recognizing objects and providing accurate positional information within images. As illustrated in Fig. 1, the model learns from labeled support samples of known categories during the training phase. In the testing phase, the model demonstrates a remarkable ability to generalize to new categories, significantly enhancing its overall adaptability and performance.
@@
Figure 1: Demonstration of the Few-Shot Object Localization (FSOL) task. During the training phase, the model predicts the location map based on given support samples and their corresponding query image. It then adjusts its parameters by by minimizing the Mean Squared Error loss between the ground truth and the predicted location map. In the testing phase, the trained model predicts the location map of novel class samples on corresponding query images that were not appear in the training phase.
@@
This study aims to advance the research field by introducing a high-performance benchmark model for the FSOL task. Within this task, we have identified two primary challenges: 1) Appearance gap between intra-class objects: Significant variations among objects within the same class in the query image create an appearance gap compared to the support image samples, impacting query accuracy (see Fig. 2(a)). 2) Object omission due to inter-object occlusion: The model struggles to accurately distinguish dense, overlapping objects in the query image, leading to a decrease in localization recall rate (see Fig. 2(b)).
To tackle the key challenges in the FSOL task, we design a dual-path feature augmentation module. To handle intra-class variations in shape, size, and orientation, we use a deformable convolution branch [13, 14], which enhances localization performance by adapting to feature variations. To reduce object omission, we implement a cross-center difference convolution branch [15], improving feature discriminability by capturing gradient difference. Additionally, we introduce 3D convolution to capture image structure, texture, and patterns, enhancing feature representation and model performance. The query image performs 3D convolution with the support image to generate a similarity map reflecting object locations.
Further, leveraging the original query image to enhance the obtained similarity map has emerged as a promising strategy. Current methods often add the query feature directly to the similarity feature map [16], using a residual connection technique to retain object information from the original image and optimize the similarity feature map. However, this direct addition strategy introduces considerable noise from the query image, making it unsuitable for localization tasks that demand high accuracy. Therefore, inspired by self-support matching [17], we utilize the similarity matrix calculated between the query image and the similarity map for weighting. This approach aims to more accurately incorporate the query image’s information while reducing undesirable noise.
In this paper, we introduces the pioneering task of FSOL and presents an innovative high-performance benchmark. To tackle the challenges posed by significant intra-class variations and target occlusions in localization tasks, we design a dual-path feature augmentation module that aims to enhance appearance correspondence and gradient discrimination between support and query features. Moreover, to effectively leverage information in query images for enhancing similarity maps, we introduce a self query module to explore intricate associations between feature maps and query images. Experimental results demonstrate the substantial performance improvement of our approach in the FSOL task, establishing an efficient benchmark for further research in object localization under limited data scenarios. In summary, the contributions of this paper can be outlined as follows:
  • For the first time, we define the task of few-shot object localization, proposing a new research direction for object localization in scenarios with limited labeled data, and establish a high-performance benchmark.
  • The dual-path feature augmentation module is designed to simultaneously enhance the deformation and gradient association between support and query images, significantly improving localization performance.
  • The self-query module uses similarity matrix weighting to leverage the query image for enhancing the similarity map while avoiding excessive noise interference.
@@
Figure 2: Difficulties in few-shot object localization that cause negative influences: a) Appearance gap between intra-class objects; b) Object omission due to inter-object occlusion.
@@ II Related Work
This paper is the first work to explore the task of few-shot object localization. Therefore, in this section, we briefly review the research in other fields related to this paper. First, we outline some classic works for the object localization task, then briefly introduce the development of few-shot learning, and finally introduce the currently prevalent works for few-shot object counting.
@@ II-A Object Localization
Object localization aims to precisely determine the position of objects in images or videos. Unlike object detection, which involves bounding box information, object localization focuses solely on object position. This localization process typically involves two stages: generating a location map through model prediction, followed by post-processing to extract object position details [18, 19, 20]. The evolution from density maps [21, 22, 23] to location maps [24, 18, 19] emphasizes accurate object localization and high-quality gradient data for model training. Traditional density maps struggle with precise object localization, prompting innovations like the IDT map [24], designed to provide accurate object positions through distance flipping. Despite improvements, IDT maps often contain significant background noise. Addressing this, Liang et al. [18] introduced the Focal Inverse Distance Transform (FIDT) map, offering enhanced gradient information. However, issues like rapid foreground decay and slow background decline persist. To address these, Li et al. [19] proposed using exponential functions to moderate foreground decay and expedite background zeroing, resulting in more optimal location maps.
@@ II-B Few-shot Learning
Few-shot learning methods offer competitive performance even with limited annotated data, a highly promising attribute. This paper proposes integrating few-shot learning techniques with object localization tasks to explore innovative methods for improving object localization performance in few-shot scenarios.
@@ II-C Few-shot Object Counting
The Few-shot Object Counting task aims to describe object categories using support images and then count the instances of these categories in the query image. Previous research [32] primarily focused on enumerating specific objects within single categories, constraining the task’s generalization capability. Lu et al. [33] introduced class-agnostic counting, developing a GMN model capable of counting objects across any category by reframing the problem as an image self-similarity matching task. In contrast, Yang et al. [34] employed multi-scale computation to assess similarity between support and query images, transforming object counting into a similarity matching challenge. Due to the absence of dedicated few-shot object counting datasets, Ranjan et al. [35] introduced the FSC-147 dataset, which includes 147 object categories and 6000 images, and proposed the FamNet model based on similarity comparison techniques. More recently, You et al. [16] integrated feature enhancement and similarity matching concepts, achieving state-of-the-art performance in few-shot object counting. These studies contribute both theoretically and empirically to advancing the few-shot object counting task.
@@ III Method
This study proposes a novel Few-Shot Object Localization (FSOL) task to enhance object localization accuracy and generalization in few-shot scenarios. To this end, we employ three key strategies, as shown in Fig. 3. Firstly, our method emphasizes capturing deformation and gradient information within the image to better understand the association between query and support features. Secondly, we integrate 3D convolution to comprehensively consider various channels’ impact on feature similarity comparison. Lastly, we introduce a module utilizing original query features to guide similarity distribution, allowing the model to utilize optimized similarity information for identifying and locating challenging samples. This reduces object omissions and enhances the model’s generalization.
Figure 3: Demonstration of our FSOL Pipeline. Given the query and support images, the query feature $F_{Q}$ is extracted from the query image while the support feature $F_{S}$ is cropped from $F_{Q}$. The Dual-path Feature Augmentation (DFA) module first enhances deformation and gradient information in both $F_{Q}$ and $F_{S}$ through deformation and gradient branches, outputting the deformation-enhanced $F_{Q}^{D}$, $F_{S}^{D}$ as well as gradient-enhanced $F_{Q}^{C}$ and $F_{S}^{C}$. Then, DFA performs 3D convolution on the stacked $F_{Q}^{D}$ and $F_{S}^{D}$ using stacked $F_{Q}^{C}$ and $F_{S}^{C}$ as convolution kernel weights, obtaining the similarity map $S$ between query and support images. Then, the Self Query (SQ) module accepts $S$ as input and uses the original $F_{Q}$ to guide the object’s distribution information in $S$, subsequently outputting the optimized similarity map $S_{SQ}$. The $S_{SQ}$ then be sent to regression head to get the final location map.
@@ III-A Dual-path Feature Augmentation
In this study, we introduce a Dual-path Feature Augmentation (DFA) module aimed at mitigating significant intra-class appearance gaps and addressing object omission challenges in FSOL task. The DFA module comprises two key components: a deformation branch and a gradient branch. The deformation branch facilitates the model to learn the matching relationships within intra-class features, thereby improving localization accuracy by capturing deformable features in both support and query images. On the other hand, the gradient branch emphasizes comparing gradient similarities between support and query images during similarity computation. This enhances the model’s ability to discern object features, thereby effectively reducing object omissions. The subsequent sections will elaborate on the vanilla 2D convolution, followed by a detailed explanation of the deformation branch and gradient branch.
@@ III-A1 Vanilla 2D Convolution
In convolutional neural networks, the vanilla 2D convolution serves as a fundamental operator, comprising two primary steps: (i) Sampling local neighboring regions on the input feature map $x$ using $N$ sampling points; (ii) Aggregating the sampled values through learnable weights $w$. Therefore, for the position $p$ on the input feature map $x$, the corresponding value at the same position $p$ on the output feature map $y$ can be mathematically expressed as:
@@
where $p_{k}$ represents the predefined offset of the enumerated position $p$ in the local neighbor area. For example, when $N=9$, there are nine sampling points, $p_{k}=\{(-1,-1),(-1,0),\ldots,(0,1),(1,1)\}$. In this paper, we initially apply vanilla 2D convolution to process the image. For an input image $\mathbb{R}^{1\times 3\times 512\times 512}$, we use the first three stage outputs of ResNet50, resulting in outputs of 256, 512, and 1024 channels, respectively. These outputs are concatenated into a feature tensor $F\in\mathbb{R}^{1\times 1792\times 128\times 128}$. Subsequently, query features $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ are obtained through the aforementioned vanilla 2D convolution. For our primary experiment, the support feature $F_{S}\in\mathbb{R}^{1\times 256\times x\times y}$ is cropped from the query feature $F_{Q}$ based on the dimensions of the original input sample. Following an adaptive pooling layer, the support feature is resized to $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$. Subsequently, these features are forwarded to both the deformation branch and the gradient branch for further processing.
@@ III-A2 Deformation Branch
In FSOL task, the object within the query image often exhibits substantial intra-class variation, distinct from the object depicted in the support image, thereby resulting in inaccurate queries. To address this issue, we aim to enhance the model’s capability to capture deformable information present in both query and support features.
To this end, we incorporate a Deformable Convolution (DC [13, 14]) branch. Unlike traditional convolution, which relies on fixed sampling points within the receptive field, DC introduces an offset to each sampling point, thereby making their positions variable. This offset is generated by another convolution layer. In Fig. 4(a), we depict the differences between sampling points in vanilla and deformable convolution. While vanilla convolution tends to integrate redundant information into features, deformable convolution diminishes redundancy by iteratively optimizing sampling points. To assess the relevance of the introduced area, DC incorporates an additional weight coefficient. When the area is deemed irrelevant, this coefficient is set to 0. Mathematically, for the same position $p$ in the input feature map $x$ and the output feature map $y$, the relationship between the values $x(p)$ and $y(p)$ after deformable convolution can be expressed as follows:
@@
where $\Delta p_{k}$ represents the offset learned by each sampling point, and $\Delta m_{k}$ represents the weight coefficient in the range [0, 1]. In our framework, both the original query and support features fed into deformation branch are processed by a shared DC layer to highlight their deformation information. Specifically, the query feature $F_{Q}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and support feature $F_{S}\in\mathbb{R}^{1\times 256\times 3\times 3}$ each go through a deformable convolution layer, resulting in the enhanced query feature $F_{Q}^{D}\in\mathbb{R}^{1\times 256\times 128\times 128}$ and the corresponding enhanced support feature $F_{S}^{D}\in\mathbb{R}^{1\times 256\times 3\times 3}$. This design not only reduces the amount of computation but also unifies the deformation distribution of both query and support features, alleviating the appearance gap between intra-class samples.
Figure 4: Demonstration of two convolutional strategies for enhancing deformation and gradient information in support and query images: (a) Deformable Conv: Vanilla convolution utilizes fixed sampling points, potentially introducing noise, while deformable convolution adjusts sampling points, reducing background noise and improving adaptability; (b) CCD-Conv: Cross-center difference convolution computes differences between neighboring pixels around the central pixel and uses these differences as weights to generate the final output. This approach captures subtle image changes like texture, edges, and fine details.